An ac generator supplies an rms (not peak) voltage of 25 V at 100.0 Hz. The generator is connected to a circuit whose capacitance is 1.50 LaTeX: \muμF. What is the rms current of the circuit?
Answer #1

Xc = 1/2pi x 100 x 1.5 x 10^-6 = 1061.03 ohm

Current, I = 25/1061.03 = 0.024 A

Comment in case any doubt please rate my answer ...